# Assessing Vulnerability in Smart Contracts: The Role of Code Complexity Metrics in Security Analysis

### v1

@@ Abstract
Codes with specific characteristics are more exposed to security vulnerabilities. Studies have revealed that codes that do not adhere to best practices are more challenging to verify and maintain, increasing the likelihood of unnoticed or unintentionally introduced vulnerabilities. Given the crucial role of smart contracts in blockchain systems, ensuring their security and conducting thorough vulnerability analysis is critical. This study investigates the use of code complexity metrics as indicators of vulnerable code in Solidity smart contracts. We highlight the significance of complexity metrics as valuable complementary features for vulnerability assessment and provide insights into the individual power of each metric. By analyzing 21 complexity metrics, we explored their interrelation, association with vulnerability, discriminative power, and mean values in vulnerable versus neutral codes. The results revealed some high correlations and potential redundancies among certain metrics, but weak correlations between each independent metric and vulnerability. Nevertheless, we found that all metrics can effectively discriminate between vulnerable and neutral codes, and most complexity metrics, except for three, exhibited higher values in vulnerable codes.
@@ I Introduction
Ethereum, referenced as [1], is a decentralized computing platform and operating system based on blockchain technology. It is an open-source platform that caters to diverse business scenarios by offering Solidity, a Turing-complete programming language, and the Ethereum Virtual Machine (EVM). Developers can utilize these tools to deploy smart contracts [2] on the Ethereum blockchain. smart contracts find widespread applications in various industries, including decentralized finance (DeFi) within financial services [3], infrastructure development [4], the internet of things (IoT) [5], the gaming industry [3], healthcare [6, 7], the metaverse [8], and numerous other sectors [3]. In contrast to traditional applications, smart contracts are executed automatically in shortcodes, which safeguards them against manipulation during the execution process. Open-source availability is a common characteristic of widely-used smart contracts, infusing trust among users regarding the integrity of the code. Since smart contracts operate on distributed nodes within the blockchain, once a contract is executed and deployed in a block, it becomes quite challenging to upgrade or modify the contract [9]. This is due to the inherent immutability of the blockchain, where tampering with any block is virtually impossible. Furthermore, smart contracts have the capability to hold funds within their associated blocks on the blockchain. These unique properties of smart contracts make them an attractive target for hackers. While the underlying blockchain technology itself is highly secure and resistant to hacking attempts, smart contracts can still be vulnerable to attacks. This vulnerability often stems from factors such as programmer negligence or errors in the code. Therefore, implementing robust vulnerability detection mechanisms becomes essential to identify and address any potential security flaws in smart contracts prior to their deployment [10, 11]. The security of smart contracts is of utmost importance, as evidenced by the significant amount of money lost in real-world smart contract vulnerabilities each month. A notable example is the attack on the Decentralized Autonomous Organization (DAO) in 2016, where 3.6 million Ether was stolen using a simple recursive vulnerability [12]. Additionally, according to blockchaingroup.io  [^1] https://blockchaingroup.io/compliance-and-regulation/top-10-crypto-losses-of-2024-hacks-frauds-and-exploits/, the top 10 crypto hacks in 2023 alone resulted in the theft of 1.06 billion US dollars by hackers. Table I provides an overview of the amounts stolen in each smart contract hack. These staggering numbers serve as evidence of the critical importance of ensuring the security of smart contracts.
@@
The rest of the paper is structured as follows. Section two presents our contribution and outlines the research questions we aim to answer. Section three provides background and related work, offering the necessary knowledge to understand this study and reviewing previous work on the topic. Section four details each complexity metric used in this study. Section five describes the dataset utilized. Section six, the experimental setup, explains the techniques and analysis methods employed to answer the research questions. Section seven presents the results of our analysis, addressing each research question in detail and analyzing the answers. Section eight, the discussion, explores the limitations of this study and suggests future directions. Lastly, section nine concludes the paper, summarizing the answers to the research questions and the overall findings.
@@ II Contribution & Research Questions
The primary focus of this study is to investigate the impact of complexity metrics on vulnerabilities present in Solidity smart contracts. While existing static analysis tools are designed to detect specific vulnerability patterns, complexity metrics offer a complementary approach to identifying vulnerable areas within smart contracts. Our specific research objectives are as follows:
@@ III Background & Related Work
Researchers have been actively proposing various methods and tools to identify security vulnerabilities in smart contracts. These approaches include static and dynamic code analysis, as well as formal methods [13] and Machine learning.
@@ III-A Static Analysis
Static code analysis involves examining the source code of a smart contract without executing it, aiming to detect potential vulnerabilities through code inspection. Slither [14] is a Python 3-based open-source static analysis framework that relies on the Solidity compiler (Solc) as its dependency. It offers swift, reliable, precise, and comprehensive information about smart contracts. Slither makes use of SlithIR, an intermediate representation designed for practical analysis of Solidity code. In addition to bug detection, it provides suggestions for code optimizations and enhances the comprehension of smart contract code [15]. In addition to Slither, there are several other noteworthy static analysis tools for smart contracts, such as SolAnalyser [16] and EtherTrust [17].
III-B Dynamic Analysis
Dynamic code analysis, on the other hand, involves analyzing the behavior of a smart contract during execution to identify any security flaws.
III-C Machine Learning
Machine learning techniques have proven to be valuable in the detection of software vulnerabilities [18]. By analyzing vast amounts of data and identifying patterns or anomalies, machine learning algorithms can effectively pinpoint potential weaknesses in software. This approach serves as a complementary method, augmenting other analysis approaches by providing automated analysis capabilities and enabling scalability in the vulnerability detection process. Machine learning’s ability to learn from data and identify subtle patterns makes it a promising tool for enhancing the efficiency and effectiveness of software vulnerability detection.
A recent machine learning (ML) model, referred to as MODNN [19], demonstrated success in utilizing a publicly available dataset to evaluate its approach. By leveraging this dataset, the MODNN model showcased its efficacy in analyzing and addressing vulnerable smart contracts. The proposal introduces the Multiple-Objective Detection Neural Network (MODNN) as a scalable smart contract vulnerability detection tool, leveraging the ML MOD (Multiple-Objective Detection) method. The aim is to enhance the security of cryptocurrency systems. A novel approach is also presented, focusing on extracting implicit features from operation sequences to support the scalability of the MODNN model. By combining these advancements, the research aims to provide a more efficient and effective solution for detecting vulnerabilities in smart contracts within cryptocurrency systems.
The researchers in DR-GCN [20] investigate the use of graph neural networks (GNN) and expert knowledge for detecting vulnerabilities in smart contracts. They focus on capturing the extensive control-flow and data-flow semantics of the source code by representing it as a contract graph. To enhance the interpretability of the graph, they introduce a node elimination phase to normalize it and identify critical nodes. Subsequently, they propose a unique temporal message propagation network that extracts graph features from the normalized graph. These graph features are then combined with expert patterns that have been specifically designed for vulnerability detection. The approach outlined in the paper aims to leverage both GNN and expert knowledge to improve the accuracy and effectiveness of vulnerability detection in smart contracts.
In Soliaudit, a study conducted by Liao et al. [21], machine learning and fuzz testing techniques are employed for assessing vulnerabilities in smart contracts. SoliAudit utilizes machine learning technology, specifically using Solidity machine code as learning features, to verify 13 different types of vulnerabilities. The researchers also develop a gray-box fuzz testing mechanism comprising a fuzzer contract and a simulated blockchain environment for online transaction verification. Notably, SoliAudit distinguishes itself from previous research systems by its ability to detect vulnerabilities without relying on expert knowledge or predefined patterns.
Eth2Vec [22] is a machine-learning-based static analysis tool designed to detect vulnerabilities in smart contracts. It operates by automatically learning the features of vulnerable Ethereum Virtual Machine (EVM) bytecodes using a neural network for natural language processing. This approach enables Eth2Vec to acquire tacit knowledge and expertise in identifying vulnerabilities. By leveraging the learned knowledge, Eth2Vec can effectively detect vulnerabilities in smart contracts by comparing the code similarity between the target EVM bytecodes and the bytecodes it has already learned. These successful evaluations highlight the potential of ML models to harness public datasets, enabling researchers to refine and enhance their approaches to tackle complex problems. These combined efforts contribute to the ongoing development of effective approaches for identifying and mitigating security vulnerabilities in smart contracts.
While machine learning models have made significant steps, they still face challenges in effectively detecting vulnerabilities within smart contracts. To maintain their capabilities, the incorporation of additional features is crucial. Notably, the potential of software metrics, particularly complexity metrics, has yet to be thoroughly explored in terms of their effectiveness in identifying vulnerabilities in the domain of smart contracts.
III-D Complexity and Vulnerability
@@
Defining program complexity is a crucial step in understanding software systems. According to Basili, complexity can be seen as a measure of the resources utilized by a system while interacting with a piece of software to accomplish a specific task [23]. The interpretation of complexity depends on the interacting system, whether it is a computer or a programmer. In the context of a computer, complexity is determined by the execution time and storage requirements for computation. On the other hand, when considering the interaction between a program and a programmer, complexity is characterized by the challenges involved in tasks such as coding, debugging, testing, and modifying the software. The term “software complexity” often refers to the interaction between a program and a programmer engaged in programming activities [24].
@@
Security experts assert that complexity poses a significant threat to security [25, 26]. Complex code can introduce subtle vulnerabilities that are challenging to detect and diagnose [26], thereby offering attackers greater opportunities for exploitation. Furthermore, complex code presents difficulties in comprehension, maintenance, and testing. Consequently, complex code is more prone to faults compared to simpler code. As attackers exploit these faults in a program, complex code becomes more vulnerable than its simpler counterparts. Until now, the use of complexity metrics as indicators of syntactic and semantic information extracted from smart contracts has not been extensively explored [27]. However, the effectiveness of complexity metrics in identifying vulnerable code has been extensively studied and validated in various applications and programming languages [28]. Researchers have leveraged techniques like machine learning and statistical models to demonstrate the correlation between code complexity and the presence of vulnerabilities.
@@
III-E Smart Contract Architecture
@@
Figure 2: Structure of a Solidity smart contract
@@ V Dataset and Labeling
For our research, we utilized the dataset provided by Liu et al [39]. They made their work publicly available on their GitHub repository, enabling further investigation. We use the “Resource 3” on their GitHub [^2] https://github.com/Messi-Q/Smart-Contract-Dataset, which contains 8 types of vulnerabilities. The vulnerability types they measured can be found in Table III. This newly published dataset, released in 2023, includes 2,953 smart contracts and a total of 16,239 contracts which contained 258 vulnerable and 15981 neutral contracts.
@@ VI Experimental Setup
To answer our questions, we employ several statistical techniques and machine-learning supervised classifiers.
VI-A Statistical Techniques
@@
Confidence Interval: The primary objective of computing a statistic from a random sample is to approximate the mean of the entire population. However, the accuracy of this approximation is always a concern. To address this issue, a confidence interval is employed, which provides a range of values likely to contain the true population parameter of interest. The confidence level for the interval is chosen by the specialist, typically set at 95%. This means that if the population is sampled multiple times, the resulting interval estimates would encompass the true population parameter in approximately 95% of cases [42]. The confidence interval provides a valuable tool to gauge the reliability and precision of the sample statistic in estimating the characteristics of the entire population.
VI-B Machine Learning Classifiers
To assess the predictive capability of Machine Learning models using these complexity metrics for vulnerable contract identification and answer question 5, we employed various ML classifiers along with Multilayer Perceptron. The evaluation is conducted using 10-fold cross-validation on our dataset, ensuring the robustness and reliability of the results. We report four commonly used evaluation metrics: accuracy, precision, recall, and F1-measure. In the evaluation of our classifiers, we utilize the basic criteria of True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN). These base values serve as the foundation for calculating other performance indicators.
  • A True Positive (TP) occurs when the classifier correctly identifies a positive instance (vulnerable contract) as positive. It represents a correct prediction of vulnerability.
  • A True Negative (TN) is the correct identification of a negative instance (non-vulnerable contract) as negative. It signifies a correct prediction of non-vulnerability.
  • On the other hand, a False Positive (FP) happens when the classifier incorrectly classifies a negative instance as positive (non-vulnerable contract as vulnerable).
  • Lastly, a False Negative (FN) occurs when the classifier misclassifies a positive instance as negative (vulnerable contract as non-vulnerable).
The relationship between these four criteria is summarized in Table IV.
TABLE IV: Basic evaluation criteria
Criteria
TP
Correct Result
FP
Unexpected result
FN
Missing result
TN
Correct absence of result
In our experimental results, we report several performance metrics to evaluate the effectiveness of our models. Similar to MODNN, we calculate the “weighted average” scores of these metrics to account for the class imbalance. The “weighted average” means that the scores are calculated by considering the class distribution and giving more weight to the underrepresented class (vulnerable contracts) in the evaluation.
The performance metrics we report are as follows:
  • Accuracy: It measures the fraction of correct predictions made by our models. It is calculated as the sum of true positives and true negatives divided by the total number of predictions.
    $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (2)
  • Recall: Also known as sensitivity or true positive rate, indicates how many of the vulnerable contracts our models could find. It is calculated as the ratio of true positives to the sum of true positives and false negatives.
    $Recall=\frac{TP}{TP+FN}$ (3)
  • Precision: Reports how many of the vulnerable contracts that our models could find were indeed vulnerable. It is calculated as the ratio of true positives to the sum of true positives and false positives.
    $Precision=\frac{TP}{TP+FP}$ (4)
  • F1 score: It is the harmonic mean of precision and recall, providing a balanced measure of the model’s performance. It is calculated using the formula:
    $F1=\frac{2*Precision*Recall}{Precision+Recall}$ (5)
  • FPR: The False Positive Rate and False Negative Rate are two important metrics used to evaluate the performance of binary classification models, such as ours. These rates help us understand how well a model is able to distinguish between the two classes. The False Positive Rate is the ratio of the number of false positives to the total number of actual negatives (true negatives + false positives). In this context, FPR refers to the proportion of neutral codes that are incorrectly classified as vulnerable. It is calculated by:
    $FPR=\frac{FP}{TN+FP}$ (6)
  • FNR: The False Negative Rate is the ratio of the number of false negatives to the total number of actual positives (true positives + false negatives). In this context, FNR refers to the proportion of vulnerable codes that are incorrectly classified as neutral. FNR is calculated by:
    $FPR=\frac{FN}{TP+FN}$ (7)
Some popular Machine learning classification models are used in this context. For classification, we utilize the following models:
  • K-Nearest Neighbors (KNN) with K values of 1, 3, and 5. Higher values of K are not considered to prevent underfitting.
  • Decision Tree classifier based on entropy.
  • AdaBoost classifier with 300 estimators.
  • Random Forest classifier with 100 estimators.
  • Multi-Layer Perceptron classifier
All models are implemented using Scikit-learn [43], a popular machine-learning library in Python.
VI-C Managing the Dataset
To address the data imbalance, we apply class weighting during the model-tuning process and weighted averaging while reporting the results. This ensures that the models can effectively handle the imbalanced nature of our dataset and produce more accurate results.
Alternative approaches for handling imbalanced datasets, such as SMOTE (Synthetic Minority Over-sampling Technique) [44], might still be necessary. SMOTE addresses class imbalance by combining oversampling and undersampling methods. Oversampling involves generating synthetic samples for the minority class by interpolating between existing instances, effectively increasing the number of minority class samples. However, this method is not suitable for our work, as it would generate an artificial dataset with metric values that may never be seen in real code. Conversely, undersampling involves randomly removing instances from the majority class to reduce its dominance. This method may be useful since it does not create artificial metric values. By balancing the dataset through these combined techniques, SMOTE can improve model performance on imbalanced datasets.
To provide a comprehensive evaluation, we report the average classification results over the ten folds, along with the corresponding standard deviation. This allows us to assess the consistency and stability of the models’ performance across different subsets of the dataset.
For our execution environment, we utilized Google Colab  [^3] https://colab.research.google.com/ in its free-plan hardware accessibility mode, on its CPU. Google Colab is a cloud-based platform used to execute our programs and conduct experiments. Google Colab provides a convenient environment for running Python code, leveraging its powerful computational resources and integration with popular libraries and frameworks such as Scikit-learn.

### v2

@@ Abstract
Codes with specific characteristics are more exposed to security vulnerabilities. Studies have revealed that codes that do not stick to best practices are more challenging to verify and maintain, increasing the likelihood of unnoticed or unintentionally introduced vulnerabilities. Given the crucial role of smart contracts in blockchain systems, ensuring their security and conducting thorough vulnerability analysis is critical. This study investigates the use of code complexity metrics as indicators of vulnerable code in Solidity smart contracts. We highlight the significance of complexity metrics as valuable complementary features for vulnerability assessment and provide insights into the individual power of each metric. By analyzing 21 complexity metrics, we explored their interrelation, association with vulnerability, discriminative power, and mean values in vulnerable versus neutral codes. The results revealed some high correlations and potential redundancies among certain metrics, but weak correlations between each independent metric and vulnerability. Nevertheless, we found that all metrics can effectively discriminate between vulnerable and neutral codes, and most complexity metrics, except for three, exhibited higher values in vulnerable codes.
@@ I Introduction & Motivation
Ethereum [1], is a decentralized computing platform and operating system based on blockchain technology. It is an open-source platform that caters to diverse business scenarios by offering Solidity, a Turing-complete programming language, and the Ethereum Virtual Machine (EVM). Developers can utilize these tools to deploy smart contracts [2] on the Ethereum blockchain. smart contracts find widespread applications in various industries, including decentralized finance (DeFi) within financial services [3], infrastructure development [4], the internet of things (IoT) [5], the gaming industry [3], healthcare [6, 7], the metaverse [8], and numerous other sectors [3]. In contrast to traditional applications, smart contracts are executed automatically in shortcodes, which safeguards them against manipulation during the execution process. Open-source availability is a common characteristic of widely-used smart contracts, infusing trust among users regarding the integrity of the code. Since smart contracts operate on distributed nodes within the blockchain, once a contract is executed and deployed in a block, it becomes quite challenging to upgrade or modify the contract [9]. This is due to the inherent immutability of the blockchain, where tampering with any block is virtually impossible. Furthermore, smart contracts have the capability to hold funds within their associated blocks on the blockchain. These unique properties of smart contracts make them an attractive target for hackers. While the underlying blockchain technology itself is highly secure and resistant to hacking attempts, smart contracts can still be vulnerable to attacks. This vulnerability often stems from factors such as programmer negligence or errors in the code. Therefore, implementing robust vulnerability detection mechanisms becomes essential to identify and address any potential security flaws in smart contracts prior to their deployment [10, 11]. The security of smart contracts is of utmost importance, as evidenced by the significant amount of money lost in real-world smart contract vulnerabilities each month. A notable example is the attack on the Decentralized Autonomous Organization (DAO) in 2016, where 3.6 million Ether was stolen using a simple recursive vulnerability [12]. Additionally, according to blockchaingroup.io  [^1] https://blockchaingroup.io/compliance-and-regulation/top-10-crypto-losses-of-2024-hacks-frauds-and-exploits/, the top 10 crypto hacks in 2024 alone resulted in the theft of 1.03 billion US dollars by hackers. Table I provides an overview of the amounts stolen in each smart contract hack. These staggering numbers serve as evidence of the critical importance of ensuring the security of smart contracts.
Complexity metrics are used in this study because they provide a quantifiable measure of a smart contract’s structural intricacy[13], which has been historically linked to software defects and vulnerabilities [14, 15, 16]. Their use is justified and motivated by several key reasons.
Relevance to vulnerability prediction: Complexity has long been associated with software defects in traditional programming languages. Studies have shown that as complexity increases, the likelihood of errors and vulnerabilities also rises [17]. Since smart contracts execute autonomously and handle financial assets, even minor defects can lead to catastrophic failures. By using complexity metrics, the study aims to assess whether similar patterns hold for Solidity contracts.
Objectivity and Quantifiability Complexity metrics provide an objective and reproducible way to measure code characteristics [18]. Unlike qualitative assessments (e.g., code readability or subjective security evaluations), these metrics can be systematically computed, analyzed, and compared across large datasets. This makes them a suitable choice for large-scale vulnerability detection efforts.
Practical Integration into Automated Tools Automated vulnerability detection tools rely on measurable features to assess risk [19]. Complexity metrics are easy to extract from code and can be directly integrated into static analysis tools.
By focusing on complexity metrics, this study ensures a scalable, automated, and objective approach to understanding vulnerability trends in smart contracts. While other metrics may also provide valuable insights, complexity remains a fundamental property that directly influences the security and maintainability of smart contracts.
@@
The rest of the paper is structured as follows. Section two presents our contribution and outlines the research questions we aim to answer. Section three provides background and related work, offering the necessary knowledge to understand the motivation of this study and reviewing previous work on the topic. Section four details each complexity metric used in this study. Section five describes the dataset utilized. Section six, the experimental setup, explains the techniques and analysis methods employed to answer the research questions. Section seven presents the results of our analysis, addressing each research question in detail and analyzing the answers. Section eight, the discussion, explores how this study’s findings might influence smart contract best practices or developer guidelines and suggests future directions. Section nine, the threat to validity, explains the limitations and biases of this study and how they were addressed. Lastly, section ten concludes the paper, summarizing the answers to the research questions and the overall findings.
@@ II Contribution & Research Questions
The primary focus of this study is to investigate the impact of complexity metrics on vulnerabilities present in Solidity smart contracts. While existing static analysis tools are designed to detect specific vulnerability patterns[20], complexity metrics offer a complementary approach to identifying vulnerable areas within smart contracts. Our specific research objectives are as follows:
@@ III Background & Related Work
Researchers have been actively proposing various methods and tools to identify security vulnerabilities within software.
@@ III-A Static Analysis
Static code analysis involves examining the source code of a smart contract without executing it, aiming to detect potential vulnerabilities through code inspection. Slither [21] is a Python 3-based open-source static analysis framework that relies on the Solidity compiler (Solc) as its dependency. It offers swift, reliable, precise, and comprehensive information about smart contracts. Slither makes use of SlithIR, an intermediate representation designed for practical analysis of Solidity code. In addition to bug detection, it provides suggestions for code optimizations and enhances the comprehension of smart contract code. In addition to Slither, there are several other noteworthy static analysis tools for smart contracts, such as SolAnalyser [22] and EtherTrust [23].
III-B Complexity and Vulnerability
@@
Defining program complexity is a crucial step in understanding software systems. According to Basili [24], complexity can be seen as a measure of the resources utilized by a system while interacting with a piece of software to accomplish a specific task. The interpretation of complexity depends on the interacting system, whether it is a computer or a programmer. In the context of a computer, complexity is determined by the execution time and storage requirements for computation. On the other hand, when considering the interaction between a program and a programmer, complexity is characterized by the challenges involved in tasks such as coding, debugging, testing, and modifying the software. The term “software complexity” often refers to the interaction between a program and a programmer engaged in programming activities [25].
@@
Security experts assert that complexity poses a significant threat to security [26, 27]. Complex code can introduce subtle vulnerabilities that are challenging to detect and diagnose [27], thereby offering attackers greater opportunities for exploitation. Furthermore, complex code presents difficulties in comprehension, maintenance, and testing. Consequently, complex code is more prone to faults compared to simpler code. As attackers exploit these faults in a program, complex code becomes more vulnerable than its simpler counterparts. Although software complexity metrics in smart contracts have been widely explored, their precise relationship to vulnerabilities remains an open problem [15, 13, 28]. However, the effectiveness of complexity metrics in identifying vulnerable code in traditional programs has been extensively studied and validated in various applications and programming languages [29, 30]. Researchers have leveraged techniques like machine learning and statistical models to demonstrate the correlation between code complexity and the presence of vulnerabilities.
@@
III-C Smart Contract Architecture
@@
Figure 2: General structure of a Solidity smart contract
@@ V Dataset and Labeling
For our research, we utilized the dataset provided by Liu et al [39]. This paper introduces IR-Fuzz, an automated fuzzing framework designed to detect vulnerabilities in smart contracts through invocation ordering and crucial branch revisiting.
The dataset in the paper was obtained by crawling verified contracts from Etherscan [^2] https://etherscan.io, which are real-world smart contracts deployed on the Ethereum Mainnet. During this process, 5,074 duplicate contracts were removed by comparing the hash of the contract binary code. The final dataset consists of 12,515 smart contracts with available source code.
As listed in Table III, eight types of vulnerabilities were focused on in the dataset: timestamp dependency (TP), block number dependency (BN), dangerous delegatecall (DG), Ether frozen (EF), unchecked external call (UC), reentrancy (RE), integer overflow (OF), and dangerous Ether strict equality (SE). For experimentation, all smart contracts from the dataset were deployed onto a local Ethereum test network.
For ground truth labeling, vulnerability-specific patterns were defined for each type of vulnerability to provide an initial label, followed by manual verification to confirm whether a smart contract actually contained a given vulnerability. By using these predefined patterns (e.g., keyword matching), potentially vulnerable contracts were efficiently identified while reducing the time spent labeling safe contracts.
This paper uses the publicly available “Resource 3” of the dataset on their GitHub [^3] https://github.com/Messi-Q/Smart-Contract-Dataset. This “Resource 3” dataset includes 2,953 smart contracts and a total of 16,239 individual contracts which contained 258 vulnerable and 15981 neutral contracts.
@@ VI Experimental Setup
To answer our questions, we employ several statistical techniques chosen based on the previous works on complexity metrics and the dataset.
@@
Confidence Interval: The primary objective of computing a statistic from a random sample is to approximate the mean of the entire population. However, the accuracy of this approximation is always a concern. To address this issue, a confidence interval is employed, which provides a range of values likely to contain the true population parameter of interest. The confidence level for the interval is chosen by the specialist, typically set at 95%. This means that if the population is sampled multiple times, the resulting interval estimates will encompass the true population parameter in approximately 95% of cases [42]. The confidence interval provides a valuable tool to gauge the reliability and precision of the sample statistic in estimating the characteristics of the entire population.